# MDBench: Benchmarking Data-Driven Methods for Model Discovery

### v1

@@ 1 Introduction
Data-driven equation discovery—rather than relying solely on conservation laws or physical principles—has historically played a pivotal role in scientific breakthroughs. Johannes Kepler discovered the third law of planetary motion by leveraging geometrical intuition and searching for patterns in empirically gathered data (Cranmer 2023). Similarly, Edwin Hubble identified the relationship between redshift and distance through empirical observations that, at the time, lacked theoretical explanation (Kragh 2021).
A dynamical system is a mathematical framework that describes how the state of a system evolves over time through differential equations. Many complex real-world systems arising in finance, medicine, and engineering can be modeled as nonlinear dynamical systems, e.g., in physics, they model fluid dynamics (Kleinstreuer 2018); in biology, they describe neural activity (Breakspear 2017); and in engineering, they underpin control theory (Walker 2013).
@@
    We introduce the first comprehensive testbed for model discovery, covering a diverse set of methods, including linear models; genetic programming; deep learning; and large scale pre-training approaches, and spanning both ODE and PDE systems.
@@ 2.1 Overview of Benchmarks
Several efforts have been made to benchmark MD algorithms. One of the earliest large-scale initiatives is SRBench (La Cava et al. 2021), which introduced a publicly available, continuously updated benchmark [^1] https://cavalab.org/srbench/ evaluating 14 SR methods on a variety of regression problems. A later extension, SRBench 2.0 (Aldeia et al. 2025), expanded the number of evaluated algorithms to 25 and introduced additional metrics such as energy consumption. Notably, the authors concluded that no single SR method consistently outperforms others across all metrics, and they proposed several best practices for future method development such as minimizing hyperparameters and standardizing evaluation procedures.
@@
PDEBench (Takamoto et al. 2022) is a more recent benchmark focused on PDE systems that evaluated 11 physical systems using four machine learning surrogates: U-Net (Ronneberger, Fischer, and Brox 2015), Fourier Neural Operator (Li et al. 2021), and Physics-Informed Neural Networks (PINNs) (Raissi, Perdikaris, and Karniadakis 2019). While these methods are effective for solving PDEs, they are black-box models and cannot generate symbolic equations, limiting their utility for interpretable model discovery.
Gilpin et al. (Gilpin 2021) benchmarked 131 chaotic ODE systems, framing them as time series forecasting tasks. Although their study includes symbolic regression methods, only four MD methods are tested, and no evaluation is provided on the complexity or interpretability of the discovered models.
Other benchmarks have narrower or domain-specific scopes. cp3-bench (Thing and Koksbang 2025) evaluated 12 MD algorithms on 28 datasets from cosmology and astroparticle physics. CFDBench (Luo, Chen, and Zhang 2023) focuses on computational fluid dynamics (CFD) simulations and evaluated neural operators under varying conditions, without considering symbolic recovery of governing equations. The benchmark by Otness et al. (Otness et al. 2021) includes four PDE systems, but evaluated only black-box predictive models such as convolutional neural networks and $k$-nearest neighbors.
@@
Table 1: Comparison of existing benchmarks for scientific MD. #ODE refers to the number of ODE problems, #PDE refers to the number of PDE problems, #MD refers to the number of model discovery algorithms evaluated, and Noise indicates whether robustness to noise is studied.
@@ 2.2 Overview of Model Discovery Methods
Since Koza’s early work on genetic programming (GP) for symbolic regression (Koza 1994), numerous methods have emerged for data-driven model discovery. We categorize them into four main classes.
@@
GP-based methods evolve expression trees using evolutionary operators to search for equations that best fit the data. Early implementations such as gplearn [^2] https://github.com/trevorstephens/gplearn are still widely used. More modern alternatives, including PySR (La Cava et al. 2021) and Operon (Burlacu, Kronberger, and Kommenda 2020), offer improved search efficiency and parallelism.
GP methods often struggle with convergence due to their large, unstructured search spaces. They also do not naturally learn parameters from data. Neural-guided approaches like those in (Mundhenk et al. 2021) mitigate this by incorporating learned search heuristics. Recent work further accelerates GP convergence by using large language models (LLMs) as crossover and mutation operators (Shojaee et al. 2025; Meyerson et al. 2024).
@@
Linear model-based methods represent equations as sparse linear combinations of candidate basis functions (“library”). A typical form is $f=\sum_{i=1}^{k}\beta_{i}\phi_{i}$, where $\phi_{i}$ are predefined basis terms and $\beta_{i}$ are scalar coefficients. SINDy (Brunton, Proctor, and Kutz 2016) pioneered this approach using LASSO for sparsity. PDEFIND (Rudy et al. 2017) extended it to PDEs, and WSINDy (Messenger and Bortz 2021) further improved robustness by formulating equations in weak form, avoiding direct differentiation. ESINDy (Fasel et al. 2022) introduced ensemble learning for robustness in low-data or high-noise regimes. DeepMoD (Both et al. 2021) combines neural networks with sparse regression by learning spatial features via automatic differentiation.
Bayesian variants (Yuan et al. 2023; North, Wikle, and Schliep 2025; More et al. 2023) introduce uncertainty modeling and noise-aware inference. However, all LM methods are constrained by the assumption of linearity with respect to basis functions and often lack principled ways to construct these function libraries for unknown systems.
@@
LSPT methods pretrain a single model—often a Transformer architecture (Vaswani et al. 2017)—on a large corpus of symbolic regression problems. These models learn to map input–output pairs to symbolic equations, enabling rapid inference on new data once training is complete.
Neural Symbolic Regression that Scales (NeSymReS) (Biggio et al. 2021) exemplifies this approach. It uses a Set Transformer (Lee et al. 2019) as an encoder to learn latent representations of input–output mappings from synthetic equations. A decoder then generates symbolic expressions autoregressively using beam search. After generating symbolic forms, constant placeholders are refined using post-hoc optimization with methods such as BFGS (Fletcher 2000).
To overcome limitations of this two-stage design, Kamienny et al. (Kamienny et al. 2022) introduced an end-to-end approach where the Transformer directly generates complete mathematical expressions, including constants. ODEFormer (d’Ascoli et al. 2024) further extends this framework to dynamical systems, generating full differential equations from a single observed trajectory.
The main advantage of LSPT methods is their inference speed. Once trained, they can quickly generate symbolic equations for new datasets without the need for retraining or optimization, which makes them suitable for real-time applications. However, they may perform poorly on real-world systems whose structure or dynamics differ from the synthetic equations seen during pretraining (Kamienny et al. 2022).
@@
DL-based methods use neural networks to learn symbolic equations from data. One example is Equation Learner Networks (EQL) (Martius and Lampert 2016; Sahoo, Lampert, and Martius 2018), which are fully differentiable feed-forward networks that incorporate symbolic operators (e.g., sine, cosine, multiplication) as activation functions. This design enables the integration of symbolic structure into neural models and allows EQL to be combined with other architectures for scientific discovery tasks (Kim et al. 2020).
Another class of DL methods uses recurrent neural networks (RNNs) to learn a probability distribution over symbolic expressions conditioned on data. Deep Symbolic Regression (DSR) (Petersen et al. 2021) follows this approach. It trains an RNN using a risk-seeking policy gradient algorithm, where the reward is based on the predictive accuracy of sampled equations. However, the learning signal in this setup comes only from the scalar reward based on the $(X,y)$ fit, which can be weak and indirect.
To overcome this limitation, uDSR (Landajuela et al. 2022) integrates optimization techniques from genetic programming (GP), linear models (LM), and large-scale pretraining (LSPT) as inner loops within the DSR framework. This hybrid approach achieves state-of-the-art performance on the SRBench benchmark (La Cava et al. 2021).
@@ 3.1 Datasets
ODEs describe the evolution of a system’s state variables over time. Formally, a dynamical system with $d$ state variables $\mathbf{u}\in\mathbb{R}^{d}$ is governed by equations of the form:
@@
where $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Each dataset consists of observed trajectories $\mathbf{U}\in\mathbb{R}^{N_{t}\times d}$.
We adopt the ODEBench dataset (d’Ascoli et al. 2024), which includes 63 systems from a textbook by Strogatz (Strogatz 2024) and several sourced from Wikipedia. These systems span one to four state variables and cover a range of real-world phenomena. For all methods, the input consists of state variables $\mathbf{u}$, and the targets are their time derivatives $\dot{\mathbf{u}}$.
@@
PDEs describe spatiotemporal systems where each state variable depends on both space and time. A PDE system with $d$ state variables $\mathbf{u}\in\mathbb{R}^{d}$ is represented on a $D$-dimensional spatial grid $\mathbf{X}\in\mathbb{R}^{N_{x_{1}}\times\cdots\times N_{x_{D}}}$. The spatial grid is constructed from uniformly-spaced intervals along each spatial coordinate, with spacing that may differ between them (i.e., $\Delta x_{1}\neq\cdots\neq\Delta x_{D})$. The time domain $t\in[0,T]$ is uniformly spaced.
@@
Our benchmark includes both widely-studied PDE systems in the MD literature and seven systems simulating complex physical processes in fluid dynamics. Previously established datasets include Advection and Burgers (Takamoto et al. 2022); Korteweg–de Vries (KdV), Kuramoto Sivishinky (KS), Nonlinear Schrödinger (NLS) , and Reaction Diffusion (RD) (Brunton, Proctor, and Kutz 2016); Advection Diffusion (AD) (Both et al. 2021).
We observe that commonly used PDE systems in the MD literature often fail to capture the complexity of real-world phenomena, which may involve space-dependent or piecewise forcing functions, or exhibit high dimensionality. Therefore, we include the following PDE systems to our testbed: Heat (Laser) (Abali 2016); Heat (Solar), Navier–Stokes (Channel), Navier–Stokes (Cylinder), and Reaction–Diffusion (Cylinder) (Langtangen and Logg 2017).
@@ 3.2 Experimental Setup
In order to assess the performance of the discovered equations, we investigate two metrics: 1) NMSE refers to the accuracy of the predicted time derivatives defined as the Normalized Mean Square Error of the derivatives, $\text{NMSE}(\mathbf{u}_{t},\mathbf{\hat{\mathbf{u}}}_{t})=\frac{\sum_{i=1}^{N}\left(u^{(i)}_{t}-\hat{u}^{(i)}_{t}\right)^{2}}{\sum_{i=1}^{N}\left(u^{(i)}_{t}\right)^{2}+\epsilon},$ where $\epsilon$ is a small regularization constant set to $10^{-10}$. NMSE measures trajectory-level predictive fidelity. A lower NMSE indicates better agreement between the model and data, and serves as a proxy for how well the discovered dynamics reproduce observed behavior. 2) Complexity refers to the total number of nodes, constant terms, and operations in the expression tree of the equations. This metric encourages parsimony and penalizes unnecessary long or redundant terms. The SymPy package is used to parse the discovered equations and compute their complexities (Meurer et al. 2017).
@@
Since MD methods are sensitive to hyperparameters, we adopt a unified and automated tuning protocol. For each method–dataset pair, we evaluate all combinations from a predefined hyperparameter grid and select the best configuration using a composite fitness score that balances between the complexity and accuracy of the equations. Similar to (Merler et al. 2024; Shojaee et al. 2023), the fitness function is defined as
@@
Most generic MD methods were originally designed for standard supervised learning tasks, where both input features and target outputs are explicitly available in the datasets. However, in the context of PDE-based dynamical systems, the right-hand side of the equations—typically involving spatial derivatives—must be constructed from the observed data. To enable the application of these methods to PDE discovery, we preprocess each dataset to compute the required symbolic features.
Specifically, for each PDE system with $d$ state variables, we generate a symbol set that includes the state variables themselves along with their spatial derivatives up to fourth order. These are computed using second-order accurate finite difference approximations via the findiff package (Baer 2018). The resulting symbolic feature set is $\left\{u_{1},u_{2},...,u_{d}\right\}\cup\left\{\frac{\partial^{j}u_{i}}{\partial x_{k}^{j}}\mid i=1,\ldots,d;j=1,\ldots,4;k=1,\ldots,D\right\}$. The target variables, i.e., the time derivatives $\partial u_{i}/\partial t$, are also computed using the same finite difference scheme.
@@ 3.3 Methods
From the algorithm categories presented in Section 2, we selected a representative set of model discovery methods to benchmark. Selection criteria included strong reported performance in prior work and the availability of well-documented, easily adoptable implementations. The studied method include PDEFIND (Rudy et al. 2017), WSINDy (Messenger and Bortz 2021), ESINDy/EWSINDy (Fasel et al. 2022), Bayesian (More et al. 2023), DeepMoD (Both et al. 2021), SINDy (Brunton, Proctor, and Kutz 2016), EQL (Sahoo, Lampert, and Martius 2018), uDSR (Landajuela et al. 2022), PySR (Cranmer 2023), Operon (Burlacu, Kronberger, and Kommenda 2020), ODEFormer (d’Ascoli et al. 2024), and End2end (Kamienny et al. 2022). Table 2 provides an overview of the selected algorithms, including their methodological type, the systems they are applied to, and a brief description.

### v2

@@ Abstract
Codehttps://github.com/gryaklab/mdbench
Datasetshttps://zenodo.org/records/17611099
Extended versionhttps://arxiv.org/abs/2509.20529
@@ 1 Introduction
Data-driven equation discovery—rather than relying solely on conservation laws or physical principles—has historically played a pivotal role in scientific breakthroughs. Johannes Kepler discovered the third law of planetary motion by leveraging geometrical intuition and searching for patterns in empirically gathered data (cranmer2023interpretable). Similarly, Edwin Hubble identified the relationship between redshift and distance through empirical observations that, at the time, lacked theoretical explanation (kragh2021cosmology).
A dynamical system is a mathematical framework that describes how the state of a system evolves over time through differential equations. Many complex real-world systems arising in finance, medicine, and engineering can be modeled as nonlinear dynamical systems, e.g., in physics, they model fluid dynamics (kleinstreuer2018modern); in biology, they describe neural activity (breakspear2017dynamic); and in engineering, they underpin control theory (walker2013dynamical).
@@
    We introduce the first comprehensive testbed for model discovery, covering a diverse set of methods, including linear models; genetic programming; deep learning; and large scale pretraining approaches, and spanning both ODE and PDE systems.
@@ 2.1 Overview of Benchmarks
Several efforts have been made to benchmark MD algorithms. One of the earliest large-scale initiatives is SRBench (la2021contemporary), which introduced a publicly available benchmark evaluating 14 SR methods on a variety of regression problems. A later extension, SRBench 2.0 (aldeia2025call), expanded the number of evaluated algorithms to 25 and introduced additional metrics such as energy consumption. Notably, the authors concluded that no single SR method consistently outperforms others across all metrics, and they proposed several best practices for future method development such as minimizing hyperparameters and standardizing evaluation procedures.
@@
PDEBench (takamoto2022pdebench) is a more recent benchmark focused on PDE systems that evaluated 11 physical systems using four machine learning surrogates: U-Net (ronneberger2015u), Fourier Neural Operator (li2021fourier), and Physics-Informed Neural Networks (PINNs) (raissi2019physics). While these methods are effective for solving PDEs, they are black-box models and cannot generate symbolic equations, limiting their utility for interpretable model discovery.
Gilpin et al. (gilpin2021chaos) benchmarked 131 chaotic ODE systems, framing them as time series forecasting tasks. Although their study includes symbolic regression methods, only four MD methods are tested, and no evaluation is provided on the complexity or interpretability of the discovered models.
Other benchmarks have narrower or domain-specific scopes. cp3-bench (thing2025cp3) evaluated 12 MD algorithms on 28 datasets from cosmology and astroparticle physics. CFDBench (luo2023cfdbench) focuses on computational fluid dynamics (CFD) simulations and evaluated neural operators under varying conditions, without considering symbolic recovery of governing equations. The benchmark by Otness et al. (otness2021extensible) includes four PDE systems, but evaluated only black-box predictive models such as convolutional neural networks and $k$-nearest neighbors.
@@
Table 1: Comparison of existing benchmarks for scientific MD. #ODE and #PDE refer to the number of ODE and PDE problems, #MD refers to the number of model discovery algorithms evaluated, and Noise indicates whether robustness to noise is studied.
@@ 2.2 Overview of Model Discovery Methods
Since Koza’s early work on genetic programming (GP) for symbolic regression (koza1994genetic), numerous methods have emerged for data-driven model discovery. We categorize them into four main classes.
@@
GP-based methods evolve expression tree of equations using evolutionary operators to search for equations that best fit the data. Modern and efficient implementations, such as PySR (la2021contemporary) and Operon (burlacu2020operon), are widely used in the literature.
GP methods often struggle with convergence due to their large, unstructured search spaces. They also do not naturally learn parameters from data. Neural-guided approaches like those in (mundhenk2021symbolic) mitigate this by incorporating learned search heuristics. Recent work further accelerates GP convergence by using large language models (LLMs) as crossover and mutation operators (shojaee2025llm; meyerson2024language).
@@
Linear model-based methods represent equations as sparse linear combinations of candidate basis functions (“library”). A typical form is $f=\sum_{i=1}^{k}\beta_{i}\phi_{i}$, where $\phi_{i}$ are predefined basis terms and $\beta_{i}$ are scalar coefficients. SINDy (brunton2016discovering) pioneered this approach using LASSO for sparsity. PDEFIND (rudy2017data) extended it to PDEs, and WSINDy (messenger2021weak) further improved robustness by formulating equations in weak form, avoiding direct differentiation. ESINDy (fasel2022ensemble) introduced ensemble learning for robustness in low-data or high-noise regimes. DeepMoD (both2021deepmod) combines neural networks with sparse regression by learning spatial features via automatic differentiation.
Bayesian variants (yuan2023machine; north2025bayesian; more2023bayesian) introduce uncertainty modeling and noise-aware inference. However, all LM methods are constrained by the assumption of linearity with respect to basis functions and often lack principled ways to construct these function libraries for unknown systems.
@@
LSPT methods pretrain a single model—often a Transformer architecture (vaswani2017attention)—on a large corpus of symbolic regression problems. These models learn to map input–output pairs to symbolic equations, enabling rapid inference on new data once training is complete.
Neural Symbolic Regression that Scales (NeSymReS) (biggio2021neural) exemplifies this approach. It uses a Set Transformer (pmlr-v97-lee19d) as an encoder to learn latent representations of input–output mappings from synthetic equations. A decoder then generates symbolic expressions autoregressively using beam search. After generating symbolic forms, constant placeholders are refined using post-hoc optimization with methods such as BFGS (fletcher2000practical).
To overcome limitations of this two-stage design, Kamienny et al. (kamienny2022end) introduced an end-to-end approach where the Transformer directly generates complete mathematical expressions, including constants. ODEFormer (d2024odeformer) further extends this framework to dynamical systems, generating full differential equations from a single observed trajectory.
The main advantage of LSPT methods is their inference speed. Once trained, they can quickly generate symbolic equations for new datasets without the need for retraining or optimization, which makes them suitable for real-time applications. However, they may perform poorly on real-world systems whose dynamics differ from the synthetic equations seen during pretraining (kamienny2022end).
@@
DL-based methods use neural networks to learn symbolic equations from data. One example is Equation Learner Networks (EQL) (martius2016extrapolation; sahoo2018learning), which are fully differentiable feed-forward networks that incorporate symbolic operators (e.g., sine, cosine, multiplication) as activation functions. This design enables the integration of symbolic structure into neural models and allows EQL to be combined with other architectures for scientific discovery tasks (kim2020integration).
Another class of DL methods uses recurrent neural networks (RNNs) to learn a probability distribution over symbolic expressions conditioned on data. Deep Symbolic Regression (DSR) (petersen2021deep) follows this approach. It trains an RNN using a risk-seeking policy gradient algorithm, where the reward is based on the predictive accuracy of sampled equations. However, the learning signal in this setup comes only from the scalar reward based on the $(X,y)$ fit, which can be weak and indirect.
To overcome this limitation, uDSR (landajuela2022unified) extends DSR by integrating optimization techniques from GP, LM, and LSPT approaches as inner loops. This hybrid approach achieves state-of-the-art performance on the SRBench benchmark (la2021contemporary).
@@ 3.1 Datasets
ODEs describe the evolution of a system’s state variables over time. Formally, a dynamical system with $d$ state variables $\mathbf{u}=(u_{1},\ldots,u_{d})\in\mathbb{R}^{d}$ is governed by equations of the form:
@@
where $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Each dataset consists of observed trajectories $\mathbf{U}\in\mathbb{R}^{N_{t}\times d}$, where $N_{t}$ is the number of samples across time.
We adopt the ODEBench dataset (d2024odeformer), which includes 63 systems from a textbook by Strogatz (strogatz2024nonlinear) and several sourced from Wikipedia. These systems span one to four state variables and cover a range of real-world phenomena. For all methods, the input consists of state variables $\mathbf{u}$, and the targets are their time derivatives $\dot{\mathbf{u}}$.
@@
PDEs describe spatiotemporal systems where each state variable depends on both space and time. A PDE system with $d$ state variables $\mathbf{u}=(u_{1},\ldots,u_{d})\in\mathbb{R}^{d}$ is represented on a $D$-dimensional spatial grid $\mathbf{X}\in\mathbb{R}^{N_{x_{1}}\times\cdots\times N_{x_{D}}}$. The spatial grid is constructed from uniformly-spaced intervals along each spatial coordinate, with spacing that may differ between them (i.e., $\Delta x_{1}\neq\cdots\neq\Delta x_{D})$. The time domain $t\in[0,T]$ is uniformly spaced.
@@
Our benchmark includes both widely-studied PDE systems in the MD literature and seven systems simulating complex physical processes in fluid dynamics. Previously established datasets include Advection and Burgers (takamoto2022pdebench); Korteweg-de Vries (KdV), Kuramoto-Sivishinky (KS), Nonlinear Schrödinger (NLS) , and Reaction-Diffusion (RD) (brunton2016discovering); Advection-Diffusion (AD) (both2021deepmod).
We observe that commonly used PDE systems in the MD literature often fail to capture the complexity of real-world phenomena, which may involve space-dependent or piecewise forcing functions, or exhibit high dimensionality. Therefore, we include the following PDE systems in our testbed: Heat (Laser) (abali2016computational); Heat (Solar), Navier-Stokes (Channel), Navier-Stokes (Cylinder), and Reaction-Diffusion (Cylinder) (langtangen2017solving).
@@ 3.2 Experimental Setup
In order to assess the performance of the discovered equations, we investigate two metrics: 1) NMSE refers to the accuracy of the predicted time derivatives defined as the Normalized Mean Square Error of the derivatives, $\text{NMSE}(\mathbf{u}_{t},\mathbf{\hat{\mathbf{u}}}_{t})=\frac{\sum_{i=1}^{N}\left(u^{(i)}_{t}-\hat{u}^{(i)}_{t}\right)^{2}}{\sum_{i=1}^{N}\left(u^{(i)}_{t}\right)^{2}+\epsilon},$ where $\epsilon$ is a small regularization constant set to $10^{-10}$. NMSE measures trajectory-level predictive fidelity. A lower NMSE indicates better agreement between the model and data, and serves as a proxy for how well the discovered dynamics reproduce observed behavior. 2) Complexity refers to the total number of nodes, constant terms, and operations in the expression tree of the equations. This metric encourages parsimony and penalizes unnecessary long or redundant terms. The SymPy package is used to parse the discovered equations and compute their complexities (sympy).
@@
Since MD methods are sensitive to hyperparameters, we adopt a unified and automated tuning protocol. For each method-dataset pair, we evaluate all combinations from a predefined hyperparameter grid and select the best configuration using a composite fitness score that balances between the complexity and accuracy of the equations. Similar to (merler2024context; shojaee2023transformer), the fitness function is defined as
@@
Most generic MD methods were originally designed for standard supervised learning tasks, where both input features and target outputs are explicitly available in the datasets. However, in the context of PDE-based dynamical systems, the right-hand side of the equations, which typically involve spatial derivatives, must be constructed from the observed data. To enable the application of these methods to PDE discovery, we preprocess each dataset to compute the required symbolic features.
Specifically, for each PDE system with $d$ state variables, we generate a symbol set that includes the state variables themselves along with their spatial derivatives up to fourth order. These are computed using second-order accurate finite difference approximations via the findiff package (findiff). The resulting symbolic feature set is $\left\{u_{1},u_{2},...,u_{d}\right\}\cup\left\{\frac{\partial^{j}u_{i}}{\partial x_{k}^{j}}\mid i=1,\ldots,d;j=1,\ldots,4;k=1,\ldots,D\right\}$. The target variables, i.e., the time derivatives $\partial u_{i}/\partial t$, are also computed using the same finite difference scheme.
@@ 3.3 Methods
From the algorithm categories presented in Section 2, we selected a representative set of model discovery methods to benchmark. Selection criteria included strong reported performance in prior work and the availability of well-documented, easily adoptable implementations. The studied method include PDEFIND (rudy2017data), WSINDy (messenger2021weak), ESINDy/EWSINDy (fasel2022ensemble), Bayesian (more2023bayesian), DeepMoD (both2021deepmod), SINDy (brunton2016discovering), EQL (sahoo2018learning), uDSR (landajuela2022unified), PySR (cranmer2023interpretable), Operon (burlacu2020operon), ODEFormer (d2024odeformer), and End2end (kamienny2022end). Table 2 provides an overview of the selected algorithms, including their methodological type, the systems they are applied to, and a brief description.